# MEMIT-Merge: Addressing MEMIT’s Key-Value Conflicts in Same-Subject Batch Editing for LLMs

### v2

@@ Abstract
As large language models continue to scale up, knowledge editing techniques that modify models’ internal knowledge without full retraining have gained significant attention. MEMIT, a prominent batch editing algorithm, stands out for its capability to perform mass knowledge modifications. However, we uncover a critical limitation that MEMIT’s editing efficacy significantly deteriorates when processing batches containing multiple edits sharing the same subject. Our analysis reveals that the root cause lies in MEMIT’s key value modeling framework: When multiple facts with the same subject in a batch are modeled through MEMIT’s key value mechanism, identical keys (derived from the shared subject) are forced to represent different values (corresponding to different knowledge), resulting in updates conflicts during editing. Addressing this issue, we propose MEMIT-Merge, an enhanced approach that merges value computation processes for facts sharing the same subject, effectively resolving the performance degradation in same-subject batch editing scenarios. Experimental results demonstrate that when MEMIT’s edit success rate drops to around 50% at larger batch sizes, MEMIT-Merge maintains a success rate exceeding 90%, showcasing remarkable robustness to subject entity collisions.
@@
Zilu Dong  and Xiangqing Shen  and Rui Xia School of Computer Science and Engineering Nanjing University of Science and Technology, China {zldong, xiangqing.shen, rxia}@njust.edu.cn
@@ 1 Introduction
Knowledge editing techniques aim to update models’ internal knowledge without retraining. As large language models (LLMs) continue to scale up, the prohibitive cost of full model retraining has made knowledge editing increasingly crucial in this domain. Among prevalent editing algorithms, a class of algorithms, termed “Locate and Edit” methods by Zhang et al. (2024), operates on the fundamental assumption that specific knowledge representations are localized within particular regions of the model’s parameter space, enabling targeted modifications through precise manipulation of these identified regions. MEMIT Meng et al. (2023), one of the most prominent algorithms in this class, has gained significant attention Li et al. (2024); Fang et al. (2024); Gupta et al. (2024). MEMIT inherits the core architectural feature of ROME Meng et al. (2022), which localizes knowledge to specific layers and modifies the output linear layer of MLP modules to update knowledge. The distinctive advancement of MEMIT lies in its capability to perform batch-wise mass knowledge editing, enabling simultaneous modification of multiple knowledge instances within a single batch.
@@
However, our investigation reveals a critical limitation in MEMIT: When handling batches with multiple edits that share the same subject (such as “John Smith now plays basketball.” and “John Smith comes from England.” where both statements have the subject “John Smith”), the method will exhibit significant performance degradation. In contrast, edits with different subjects (such as “Jack Johnson now plays basketball” and “Paul Morand comes from China” with subjects ‘Jack Johnson” and “Paul Morand”, respectively) maintain stable efficacy. To systematically demonstrate this performance degradation, we constructed two contrastive datasets comprising batches with identical subjects versus fully unique subjects, named distinct-subject and same-subject, respectively. The experimental results are in Fig. 1, where the vertical axis represents efficacy (which means the editing success rate) and the horizontal axis indicates the batch size per edit. The results reveal that MEMIT maintains a high success rate as batch size increases when editing distinct-subject cases, but exhibits significant performance degradation for same-subject cases. More detailed experimental settings can be found in Sec. 5.3. Batch editing of same-subject knowledge is essential in practical applications. For instance, updating multiple attributes about a specific individual (such as occupation, nationality, and affiliations) inherently requires concurrent editing of same-subject facts.
@@
Our analysis reveals that the root cause lies in MEMIT’s key-value modeling framework: when handling edits involving the same subject, identical keys (derived from shared subject representations) are forced to map to divergent values (corresponding to distinct knowledge updates). This results in inconsistencies during batch editing. Specifically, MEMIT treats each knowledge instance as a key-value pair, where the input vector to the MLP’s output linear layer serves as the key, and the desired output becomes the value.  [^1] Note that the key-value here refers to the hidden state and output within the MLP module as described by Meng et al. (2022), rather than the query, key and value in the attention module. The algorithm modifies the MLP layer by computing target values based on new knowledge objects and updating weights to map existing keys to these new values. However, MEMIT’s per-fact value computation creates inherent conflicts: within a single batch, multiple entries sharing identical keys (from same-subject cases) require different value mappings. As MLP output linear layers cannot produce multiple outputs for identical inputs, these conflicting updates lead to the performance degradation. Furthermore, we observe that the closer the keys are to each other, the more significant the performance degradation.
To resolve this fundamental conflict, we propose MEMIT-Merge, an enhanced variant of MEMIT. The core insight stems from addressing MEMIT’s critical flaw: independent value computation for same-key entries. Our solution enforces value consistency by merging multiple knowledge entries sharing identical keys, i.e, the same-subject cases into a unified value representation through constrained optimization. To validate the effectiveness of our approach, we conducted experiments using the aforementioned same-subject and distinct-subject datasets. The results demonstrate that on same-subject data, MEMIT-Merge consistently outperforms the standard MEMIT as the batch size increases. Notably, when MEMIT’s success rate drops to around 50% at larger batch sizes, MEMIT-Merge maintains a success rate exceeding 90%. While on distinct-subject data, both methods exhibit comparable performance, with no statistically significant differences observed.
2 Related Work
Knowledge editing techniques for large language models (LLMs) primarily fall into two paradigms: non-parametric approaches that preserve original parameters and parametric methods that directly modify model weights. Parametric approaches, while effective for targeted updates, often introduce uncontrolled parameter perturbations that adversely affect unrelated knowledge — a challenge addressed through various constraint mechanisms. The parametric category features two dominant subclasses: One is “Meta-Learning Based Methods”, such as MEND Mitchell et al. (2022) and MALMEN Tan et al. (2024) which train meta-networks using carefully designed datasets containing both unrelated knowledge samples and paraphrased sentences, aiming to enhance generalization while minimizing collateral damage. Another is Locate-and-Edit Methods, which includes techniques such as Knowledge Neuron (KN) Dai et al. (2022), identify critical knowledge storage locations before executing precise edits. ROME Meng et al. (2022) extends this by incorporating knowledge preservation terms in its optimization objective to maintain model integrity.
Our work builds upon MEMIT, a state-of-the-art locate-and-edit approach that enables batch knowledge editing through MLP layer modifications. Building on MEMIT, many recent methods have made modifications to parameter update methods during editing or to the architecture and location of the edits. PMET Li et al. (2024) incorporates the output of the attention layer in the calculation of parameter updates. AlphaEdit Fang et al. (2024) improves upon MEMIT’s parameter matrix update method by projecting the update matrix into the null space of the original knowledge to mitigate interference with unrelated knowledge. UNKE Deng et al. (2024) extends structured knowledge editing to unstructured editing.
@@ 3.1 Preliminaries
The MEMIT framework posits that factual knowledge is stored within the MLP layer parameters in transformer-based models. Each MLP layer in the standard architecture constitutes a two-layer neural network comprising input and output linear layers, whose parameter matrices are denoted as $W_{in}$ and $W_{out}$, respectively. MEMIT refers to the hidden state in the middle of this two-layer neural network as the key, and the final output of the MLP as the value. The output linear layer can thus be regarded as a key-value mapping, which is the optimization target of the MEMIT method, corresponding to the parameter matrix $W_{out}$. Typically, during MEMIT editing, all knowledge to be edited can be placed in a single batch, allowing for the simultaneous editing of a large amount of knowledge.
MEMIT employs a triple $(s,r,o)$ to represent the subject, relation, and object of a piece of knowledge to be edited, and constructs a complete sentence based on these elements for the editing process. During the editing process of MEMIT, the key is determined by the subject of the sentence and the part preceding the subject. The value, on the other hand, needs to be a value that enables the model to output the object. Therefore, it is optimized backward from the object of the sentence:
$v=\arg\min_{v}(-\log P_{v}[o|(s,r)]).$ (1)
This process generates a $(k,v)$ pair representing the knowledge triplet $(s,r,o)$. The model subsequently employs the following formula to update $W_{out}$ to enable the key to map to the corresponding value:
@@
where $K$ and $V$ represent the sentences composed of the k and v values in a batch, respectively. $W_{0}$ denotes the original $W_{o}ut$ weights before editing, and $C$ is a constant representing the existing knowledge. In this way, the model can output the edited $o$ with $(s,r)$.
@@ 3.2 Same subject issue in MEMIT
Normally, MEMIT is capable of maintaining its efficacy without a pronounced decline in performance when the edit batch size approaches 1,000. Moreover, the edit success rate remains relatively stable even when editing up to 10,000 knowledge triples. However, we have identified a notable issue: when the edit batch encompasses knowledge triples sharing the same subject, the editing capacity of MEMIT experiences a substantial degradation.
To verify this phenomenon, we constructed two counterfactual editing datasets. In the first dataset, the subjects of the knowledge triples are all distinct. In the second dataset, the subjects of the knowledge triples are replaced by a single, fixed subject, while all other parts of these two datasets remain identical. The details of the datasets construction are provided in App. A.
@@ 4.1 Cause Analysis
In the standard MEMIT, a piece of knowledge to be edited can be represented by a knowledge triplet (subject, relation, object), and a complete sentence is constructed based on this triplet for the editing process. In this paper, we use the format “subject’s relation is object” to construct the sentence. For example, the knowledge triple (John,father,Bob) is formulated into the sentence “John’s father is Bob.”.
@@
$AKD^{(l)}=\frac{1}{\binom{|B|}{2}}\sum_{\begin{subarray}{c}e_{1}\in B\\ e_{2}\in B\end{subarray}}||k^{(l)}_{e_{1}}-k^{(l)}_{e_{2}}||_{2}$
We compute $AKD$ for all layers of the model at the subject’s last token position. As the degree of subject variation increases across sentences, the $AKD$ value proportionally rises. Conversely, when all sentences share identical subjects, the $AKD$ value remains constant at 0.
We construct sentence batches using predefined templates, where batches sharing the same template exhibited similar $AKD$ values, while distinct templates yielded significantly different $AKD$ measurements. The specific templates and the corresponding values of $AKD$ are detailed in App. B. For experimental validation, we select three $AKD$ groups (0, 10, 25) and conduct editing tests using Qwen2.5-1.5B-Instruct. As shown in Fig. 3, where $AKD$ values are computed using keys from MEMIT’s final editing layer, the results demonstrate an inverse relationship: lower $AKD$ values correspond to reduced editing success rates. This pattern remains consistent across other $AKD$ values, establishing a statistically significant negative correlation between $AKD$ and editing efficacy.
Figure 3: Datasets with different $AKD$ values and the results of edit efficacy. The lower the $AKD$ value, the more severe the decline in edit capability.
@@ 4.2 The MEMIT-Merge Approach
To address the issue of a single key corresponding to multiple distinct values, we organize all the knowledge triples to be edited within a batch into multiple “same-subject groups” based on whether their subjects are identical. Within each group, the subjects of the knowledge triple to be edited are the same.
Compared with the optimization objective of standard MEMIT in Eq. 1, our improved version aims to optimize the knowledge triples in the same-subject group to a single value
$v=\arg\min_{v}\sum_{(s,r_{j},o_{j})\in S}-\log P_{v}[o_{j}|(s,r_{j})]$
where $S$ represents the set of knowledge triples with the subject $s$, $v$ is the value to be optimized in a backward manner, and $P_{v}$ denotes the model when the value is equal to $v$.
This approach ensures that a same-subject group gets the same value while sharing the same key, thereby significantly alleviating the decline in edit efficacy observed in standard MEMIT.
@@ 5.1 Dataset
As mentioned in the Introduction section, we constructed two counterfactual knowledge editing datasets based on Wikidata, one with the same subject and the other with distinct subjects. The dataset named “same-subject” dataset, contains 100 editing knowledge triples, all with the subject John Smith. The other dataset, named “distinct-subject”, has knowledge triples with subjects unique to each other, while the relations and objects remain consistent with those in the “same-subject” dataset. For specific details on the construction of the datasets, refer to App. A.
In terms of evaluation metrics, we refer to the metrics used by Meng et al. (2023), namely Efficacy, Paraphrase, and Specificity. Efficacy is the edit success rate, which measures the probability that the edited model produces the correct answer to the sentences used for editing. Paraphrase is measured in the same way as Efficacy, but it uses paraphrased versions of the sentences used for editing. Specificity measures the probability that facts unrelated to the edit remain consistent before and after the edit.
Although we constructed our own datasets due to the lack of specialized data on the same subject in the past, addressing same-subject issue has a practical necessity. Editing multiple attributes of an entity (e.g, updating a person’s profile) is a highly realistic demand. Furthermore, changing certain information about an entity sometimes requires a chain of changes to other information about that entity. All of these make the same subject scenarios essential for practical applications.
@@ 5.2 Experimental Setup
For baselines, we selected MEMIT and an improved version of MEMIT, PMET Li et al. (2024), as our baselines. The layers to be edited for these methods were determined based on the parameter settings in the Easyedit framework Wang et al. (2024). Specifically, we edited layers 4 to 8 for the Qwen model and layers 3 to 8 for the GPT-J model. The token position for editing was consistent with the original papers, which is selecting the subject last token. In addition to the MEMIT-based methods, we also included FT-L Zhu et al. (2020), which was used for comparison in the ROME paper, as another baseline to verify that the same-subject issue exists only in methods with the MEMIT-based architecture. For FT-L, we edited layer 21 in the GPT-J and Llama model, and layer 15 in the Qwen model.
5.3 Results when Batch Size is 100
Model Dataset Method Efficiency Parapharse Specificity
Qwen2.5-1.5B-Instruct same-subject FT 0.26 0.23 0.95
MEMIT 0.30 0.24 1.00
PMET 0.23 0.17 0.99
MEMIT-Merge 0.55 0.36 0.99
distinct-subject FT 0.23 0.21 0.99
MEMIT 1.00 0.77 0.90
PMET 0.51 0.40 0.85
MEMIT-Merge 1.00 0.77 0.90
GPT-J-6B same-subject FT 0.52 0.19 0.23
MEMIT 0.27 0.21 1.00
PMET 0.26 0.21 0.98
MEMIT-Merge 0.51 0.32 1.00
distinct-subject FT 0.47 0.28 0.22
MEMIT 1.00 0.77 0.93
PMET 0.25 0.25 0.99
MEMIT-Merge 1.00 0.77 0.93
Llama-3-8B-Instruct same-subject FT 0.67 0.47 0.27
MEMIT 0.38 0.29 0.98
PMET 0.23 0.21 0.98
MEMIT-Merge 0.71 0.44 0.98
distinct-subject FT 0.73 0.58 0.24
MEMIT 0.99 0.91 0.82
PMET 0.46 0.46 0.92
MEMIT-Merge 1.00 0.91 0.81
Qwen2.5-7B-Instruct same-subject FT 0.28 0.23 0.99
MEMIT 0.31 0.25 1.00
PMET 0.23 0.18 0.99
MEMIT-Merge 0.67 0.43 0.99
distinct-subject FT 0.23 0.22 0.98
MEMIT 0.99 0.84 0.91
PMET 0.52 0.47 0.84
MEMIT-Merge 1.00 0.86 0.90
Table 1: The complete results of the four editing methods—MEMIT, MEMIT-Merge, PMET, and FT-L—on the same-subject and distinct-subject datasets at a batch size of 100. All experimental results were obtained by re-running each editing method on our dataset.
@@ 5.3 Results when Batch Size is 100
We first compared the edit success rates of standard MEMIT, PMET, MEMIT-Merge, and FT-L on the two datasets across several models.  [^2] The results for all baselines were obtained by running the code from the Easyedit framework on our datasets.
Tab. 1 presents the complete results of several editing methods on the same subject and distinct-subject datasets at a batch size of 100. Compared with the standard MEMTI, it can be observed that our method achieves an improvement in the paraphrase metric. This improvement is likely attributable to the originally low edit success rate, and by enhancing the edit success rate, we naturally improved the paraphrase performance. Regarding specificity, it is noticeable that standard MEMIT exhibits an anomalously high specificity when the subjects are the same. However, comparing this with the results from the distinct-subject dataset reveals that specificity is actually low under normal editing conditions. Therefore, we infer that the abnormally high specificity is due to the low edit success rate, which indicates a minimal impact on the original model. In contrast, our MEMIT-Merge has a higher edit success rate, resulting in a specificity that is comparable to the distinct subject scenario and the FT method. Thus, MEMIT-Merge does not negatively affect specificity.
Compared to FT-L, it can be observed that the FT-L results do not decrease when subjects are the same. This situation is consistent with our analysis in Sec. 4.1, which indicates that the decline in editing performance under the same subject is a unique issue of the MEMIT editing method. Similarly, the results of PMET also show a decline in efficacy only when the subjects are the same. The ability of MEMIT-Merge to successfully mitigate the decline in editing performance by merging multiple edits when computing the value for the same key further supports the conclusion that key collision is the cause of this problem.
@@
Figure 4: The results of MEMIT-Merge and MEMIT methods on same-subject and distinct-subject datasets using the GPT-J-6B model, showing the changes with varying batch sizes. MEMIT-Merge is capable of significantly alleviating the decline in editing performance under the same-subject condition.

### v3

@@ Abstract
As large language models continue to scale up, knowledge editing techniques that modify models’ internal knowledge without full retraining have gained significant attention. MEMIT, a prominent batch editing algorithm, stands out for its capability to perform mass knowledge modifications. However, we uncover that MEMIT’s editing efficacy significantly deteriorates when processing batches containing multiple edits sharing the same subject. Our analysis reveals this stems from MEMIT’s key value modeling framework: identical keys (derived from the shared subject) are forced to represent different values (corresponding to different knowledge), resulting in update conflicts during editing. Addressing this issue, we propose MEMIT-Merge, an enhanced approach that merges value computation processes for facts sharing the same subject, effectively resolving the performance degradation in same-subject batch editing scenarios. Addressing this issue, we propose MEMIT-Merge, an enhanced approach that merges value computation processes for facts sharing the same subject, effectively resolving the performance degradation in same-subject batch editing scenarios. Experimental results demonstrate that when MEMIT’s edit success rate drops to around 50% at larger batch sizes, MEMIT-Merge maintains a success rate exceeding 90%, showcasing remarkable robustness to subject entity collisions. The code is available at https://github.com/NUSTM/MEMIT-Merge.
@@
Zilu Dongthanks: Equal Contribution., Xiangqing Shen11footnotemark: 1, Rui Xiathanks: Corresponding Author. School of Computer Science and Engineering, Nanjing University of Science and Technology, China {zldong, xiangqing.shen, rxia}@njust.edu.cn
@@ 1 Introduction
As large language models (LLMs) continue to scale up, the prohibitive cost of full model retraining has made knowledge editing increasingly crucial in this domain. Among prevalent editing algorithms, a class of algorithms, termed “Locate and Edit” methods by Zhang et al. (2024), enables targeted modifications through precise manipulation of specific regions. MEMIT Meng et al. (2023), one of the most prominent algorithms in this class, has gained significant attention Li et al. (2024); Fang et al. (2024); Gupta et al. (2024). It extends ROME’s architecture Meng et al. (2022) and enables the simultaneous modification of multiple knowledge instances within a single update operation.
@@
However, our investigation reveals a critical limitation in MEMIT: When handling batches with multiple edits that share the same subject (such as “John Smith now plays basketball.” and “John Smith comes from England.” share the same subject “John Smith”, while “Paul Morand comes from England” has a different subject), the method will exhibit significant performance degradation. In contrast, edits with different subjects maintain stable efficacy.
@@
To systematically demonstrate this performance degradation, we constructed two contrastive datasets comprising batches with identical subjects versus fully unique subjects, named distinct-subject and same-subject, respectively. The experimental results are in Fig. 1, where the vertical axis represents efficacy (which means the editing success rate) and the horizontal axis indicates the batch size per edit. The results reveal that MEMIT maintains a high success rate as batch size increases when editing distinct subject cases, but exhibits significant performance degradation for the same subject cases. However, same subject cases are also critical in the real-world practices (e.g., updating a person’s occupation, workplace, and employer simultaneously). More detailed experimental settings can be found in Sec. 4.3.
@@
The performance degradation stems from MEMIT’s key-value modeling paradigm: identical keys (derived from shared subject representations) map to conflicting values during same-subject batch edits. MEMIT formulates knowledge updates as MLP key-value pairs where the output linear layer’s weights are adjusted to align keys with edited values.  [^1] Note that the key-value here refers to the hidden state and output within the MLP module as described by Meng et al. (2022), rather than the query, key and value in the attention module. However, when multiple edits share subjects, their identical keys require divergent value mappings - an inherent contradiction since single-layer perceptrons cannot produce multiple outputs for identical inputs.
To resolve this fundamental conflict, we propose MEMIT-Merge, an enhanced variant of MEMIT. Our key insight is to enforce value consistency by merging multiple knowledge entries that share identical keys. Experimental results show that MEMIT-Merge consistently outperforms MEMIT on same-subject dataset, maintaining a success rate above 90%, whereas MEMIT drops to around 50%. For distinct-subject data, both methods perform comparably with no significant differences.
@@ 2.1 Preliminaries
The MEMIT framework hypothesizes that factual knowledge in models is stored within the parameters of MLP layers. Each MLP layer contains input/output linear layers with parameter matrices $W_{in}$ and $W_{out}$, where $W_{out}$ serves as the key-value mapping targeted by MEMIT editing. The key corresponds to the hidden state at the MLP’s intermediate layer while the value represents the MLP’s final output.
Knowledge is represented as triples $(s,r,o)$. During editing, complete sentences are constructed from these triples. The key is determined by the subject $s$ and its contextual prefix, while the value is obtained by inversely optimizing the object $o$:
$v=\arg\min_{v}(-\log P_{v}[o|(s,r)])$ (1)
All $(k,v)$ pairs are processed in batch to update $W_{out}$ via closed-form solution:
@@
Here, $K$ and $V$ denote batched key/value matrices, $W_{0}$ represents original parameters, and $C$ is a knowledge-preservation constant.
@@ 2.2 Same subject issue in MEMIT
Normally, MEMIT is capable of maintaining its efficacy without a pronounced decline in performance when the edit batch size approaches 1,000. However, we have identified a notable issue: when the edit batch encompasses knowledge triples sharing the same subject, the editing capacity of MEMIT experiences a substantial degradation.
To verify this phenomenon, we constructed two counterfactual editing datasets. In the first dataset, the subjects of the knowledge triples are all distinct. In the second dataset, the subjects of the knowledge triples are replaced by a single, fixed subject, while all other parts of these two datasets remain identical. The details of the construction of datasets are provided in App. A.
@@ 3.1 Cause Analysis
In the standard MEMIT, a piece of knowledge to be edited can be represented by a knowledge triple (subject, relation, object), and a complete sentence is constructed based on this triplet for the editing process. In this paper, we use the format “subject’s relation is object” to construct the sentence. For example, the knowledge triple (John,father,Bob) is formulated into the sentence “John’s father is Bob.”
@@
Furthermore, we analyzed the relationship between MEMIT editing capability and key distance within a batch, finding that closer keys lead to greater capability degradation. Due to space constraints, detailed analysis is in the App. B
@@ 3.2 The MEMIT-Merge Approach
To address this issue, we develop a new optimization objective to merge the value computation of the set of knowledge with the same key:
$v=\arg\min_{v}\sum_{(s,r_{j},o_{j})\in S}-\log P_{v}[o_{j}|(s,r_{j})],$ (3)
where $S$ represents the set of knowledge triples with the same key, $v$ is the value to be optimized in a backward manner, and $P_{v}$ denotes the model when the value is equal to $v$.  [^2] For readability, the random prefix component, though part of the original MEMIT, is omitted from this equation. Nevertheless, it is incorporated identically to MEMIT in the real implementation.
Compared with Eq. 1, this approach ensures that knowledge sharing same key gets the same value. As analyzed previously, original MEMIT encounters a contradiction: it demands that identical keys yield different values via the same $W_{out}$ matrix. Our method resolves this by merging. Therefore, our approach can significantly alleviate the decrease in edit efficacy observed in standard MEMIT as evidenced in the next section.
Moreover, our methodology demonstrates superior theoretical efficiency compared to the original MEMIT. This advantage stems from the fact that our merging strategy effectively increases the batch size during gradient descent. In contrast to MEMIT’s consistent batch size of 1 per edit, MEMIT-Merge dynamically sets its batch size to the number of edits sharing the same key.
@@ 4 Experiments
Model Dataset Method Efficacy Paraphrase Specificity
Qwen2.5-1.5B-Instruct distinct-subject FT 0.24 0.21 0.99
MEMIT 1.00 0.80 0.98
PMET 0.94 0.67 0.98
AlphaEdit 1.00 0.79 0.98
PMET-Merge 0.94 0.67 0.97
MEMIT-Merge 1.00 0.80 0.98
same-subject FT 0.24 0.22 0.99
MEMIT 0.41 0.28 1.00
PMET 0.40 0.26 0.99
AlphaEdit 0.42 0.27 0.98
PMET-Merge 0.63 0.44 0.99
MEMIT-Merge 0.95 0.48 1.00
Table 1: The complete results of the four editing methods—MEMIT, MEMIT-Merge, PMET, AlphaEdit, PMET-Merge, and FT-L—on the same-subject and distinct-subject datasets at a batch size of 10. All experimental results were obtained by re-running each editing method on our dataset.
@@ 4.1 Dataset
We constructed two Wikidata-based counterfactual knowledge editing datasets: (1) a "same-subject" set with 100 triples sharing the subject John Smith, and (2) a "distinct-subject" set with unique subjects while maintaining identical relations/objects (construction details in App. A).
In terms of evaluation metrics, we refer to the metrics used by Meng et al. (2023), namely Efficacy, Paraphrase, and Specificity. Efficacy measures the edit success rate on original sentences, paraphrase measures the success rate on paraphrased sentences. Specificity measures the probability that facts unrelated to the edit remain consistent before and after the edit.
While our datasets are novel, they address critical real-world needs. Editing multiple attributes of an entity (e.g., updating a person’s profile) is a highly realistic demand, making same-subject scenarios essential for practical applications.
@@ 4.2 Experimental Setup
For MEMIT-based baselines, we use MEMIT and an improved version of MEMIT, PMET Li et al. (2024), AlphaEdit Fang et al. (2025). In addition to the MEMIT-based methods, we also included FT-L Zhu et al. (2020), which was used for comparison in the ROME paper, as another baseline to verify that the same-subject issue exists only in methods with the MEMIT-based architecture.
4.3 Results when Batch Size is 10
@@
We first compared the edit success rates of standard MEMIT, PMET, AlphaEdit, MEMIT-Merge, PMET-Merge, and FT-L on the two datasets. [^3] The results for all baselines were obtained by running the code from the Easyedit framework on our datasets.
As shown in Tab. 1, our method outperforms standard MEMIT on the same-subject dataset with improved paraphrase accuracy, attributed to enhanced edit success rates. Notably, MEMIT’s anomalously high specificity for same-subject edits (indicating ineffective editing and a minimal impact on the original model) is corrected by our approach, achieving specificity levels comparable to FT and distinct-subject scenarios. The results of other models are detailed in App. D.
Comparing the results of MEMIT, AlphaEdit and PMET to FT-L, it can be observed that the performance drop in same-subject edits is unique to MEMIT-based methods. This phenomenon is consistent with our analysis in Sec. 3.1. By resolving key collisions through key-wise value merging, MEMIT-Merge successfully mitigates this issue, empirically confirming that key collision is the root cause of MEMIT’s limitations in same-subject cases.
The “Merge” method,introduced in Sec. 3.2, is not exclusive to MEMIT. Although PMET-Merge does not achieve the same performance as MEMIT-Merge (as illustrated in Tab. 1), it can still notably mitigate performance degradation for PMET in same-subject scenarios. Our hypothesis for this difference is that PMET modifies not only the MLP but also the attention module, a dimension we did not specifically explore.
@@
Figure 3: The results of MEMIT-Merge and MEMIT methods on same-subject and distinct-subject datasets using the Qwen2.5-1.5B-Instruct model, showing the changes with varying batch sizes. MEMIT-Merge is capable of significantly alleviating the decline in editing performance under the same-subject condition.